OpenStudy (anonymous):

Find f(2)for f(x)=x^-2 by creatiing a liimit problem using the definition of the derivative and finding the limit through a table of values. Show at least three entries in your limit table.

OpenStudy (zehanz):

I think you mean f'(2). The definition of the derivative is:\[f'(x)=\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }\](the limit of the differential quotient). So to find f'(2), you need to evaluate:\[f'(2)=\lim_{h \rightarrow 0}\frac{ f(2+h)-f(2) }{ h }=\lim_{h \rightarrow 0}\frac{ \frac{ 1 }{ (2+h)^2 }-\frac{ 1 }{ 2^2 } }{ h }=...\] Because you don't really have to evaluate the limit, yo can create a table like this: h f'(2) (approx.) 0.1 .... 0.01 .... 0.001 .... So for the first approximate value of f'(2), you set h=0.1 in the differential quotient:\[\frac{ \frac{ 1 }{ (2.1)^2 }-\frac{ 1 }{ 2^2 } }{ 0.1 }\] Do the same for h=0.01 and h=0.001. You will see a pattern, and you should be able to make a guess of what the limit (=f'(2)) really is.

OpenStudy (anonymous):

So then do you mean like this F(2) 0.2 0.02 0.002

OpenStudy (zehanz):

No, f(2) is just 1/2²=1/4 (=0.25). I think, from reading the text of your problem, you have to find f'(2), so the derivative of f for x=2. They want you to use the limit definition of the derivative. Look up what I what I wrote about that. When I do all this, and take e.g. h=0.1, then this first approximation of f'(2) is -0.2324. Now you have to do the same calculation, only with smaller and smaller values for h, to get an idea what the derivative f'(2) really is.